rssLink RSS for all categories
 
icon_blue
icon_green
icon_green
icon_orange
icon_red
icon_green
icon_green
icon_orange
icon_red
icon_red
icon_green
icon_green
icon_green
icon_orange
icon_orange
icon_red
icon_green
icon_green
icon_red
icon_red
icon_green
icon_red
icon_green
icon_blue
icon_orange
icon_green
icon_green
icon_blue
icon_green
icon_blue
icon_red
icon_green
 

FS#1659 — Lien Paris/Roubaix (1er lien)

Attached to Project— Reseau Internet et Baies
Maintenance
Entire OVH Network
CLOSED
100%
Nous avons un probleme sur le 1er lien Paris/Roubaix. Les équipements
sont en erreur. Nous avons perdu 2x10G sur 4x10G disponibles. Nous
sommes en train de chercher l'origine du probleme.

Nous tournons sur le 2ème lien en 2x10G. Tout continue à fonctionner.
Par contre à partir de 17h on risque d'avoir des problemes de saturation
des liens si le 1er lien n'est pas up. Nous allons faire le necessaire
pourque ça n'arrive pas.
Date:  Thursday, 20 September 2007, 03:11AM
Reason for closing:  Done
Comment by OVH - Thursday, 06 September 2007, 17:41PM

Nous avons eu la confirmation de notre fournisseur de fibre
qu'il y a bien une coupure entre Paris et Lille à Fontaine chaalis.
Les équipes travailent sur la resolution du probleme.

http://maps.google.com/maps?f=q&hl=en&geocode=&q=fontaine+chaalis,+france&sll=37.0625,-95.677068&sspn=66.70144,81.123047&ie=UTF8&ll=49.173371,2.68399&spn=0.013762,0.019805&t=k&z=16&om=1


Comment by OVH - Thursday, 06 September 2007, 19:17PM

L'origine du probleme: une entreprise est en train de passer des
fibres optiques entre Paris et Nord de France pour 1 opérateur tout
au long du chemin où notre fibre passe. Cette entreprise le fait via
les forages dirigés (des bouts de 1-2km au lieu des tranchés ouverts).

Le probleme est qu'ils ont touché nos fibres optiques ainsi que les
fibres d'un autre opérateur. Il y a donc une ou plusieurs coupures sur
un bout de 400-600 metres. Les équipes de maintenance sont sur place
depuis ce matin (la coupure a eu lieu à 10h27) et ont déjà localisé la
coupure de l'autre opérateur. Par contre, la coupure sur nos fibres
n'est pas encore localisée (le cable est enterré et la coupure n'est
pas nette). Il se peut qu'il y ait plusieurs coupures et dans ce cas
là il faudra changer un cable sur plusieurs kilometres. La coupure a
lieu dans les champs de la ville Ermenonville.

http://maps.google.com/maps?f=q&hl=en&geocode=&q=Ermenonville,+Oise,+Picardie,+France&sll=37.0625,-95.677068&sspn=66.70144,81.123047&ie=UTF8&ll=49.124795,2.693431&spn=0.013775,0.019805&t=k&z=16&iwloc=addr&om=1

Nous n'avons pas de delai de retablissement du service.

La connexion continue via le 2ème lien à 2x10G au lieu de 4x10G.


Comment by OVH - Thursday, 06 September 2007, 23:01PM

Notre fibre optique est abimée sur 2 endroits. La distance entre
ces 2 endroits est de 650m. Les équipes inteviennent en dur dans
la terre pour extraire les fibres et les reparer (la soudure etc).

Les opérations continuent tant dans la nuit jusqu'à ce que tout soit
reparé.


Comment by OVH - Friday, 07 September 2007, 11:31AM

Spoke with the Senior Management Team from OSP Maintenance contractor to obtain
status update;
The repair team have been working throughout the night to try to locate the addi
tional fault on the cable. OTDR traces indicate that there is a second fault (hi
gh loss) approx 630 metres from the initial break. The cable section affected tr
averses a field. The initial break was caused by a drilling machine which has pu
lled the cable resulting in these additional fractures.
The OSP civils team are on site and are currently digging a trench along the cab
le route to allow a repair to take place. There is no estimated repair time at p
resent.


Comment by OVH - Friday, 07 September 2007, 13:51PM

Sur notre fibre optique il y avait une coupure qui a été reparée. Par
contre le signal est fortement perturbé ce qui fait qu'aucun lien ne
monte (il y a trop de perte). Le forage dirigé a dû appuyer sur les
fibres et les ecraser sur une distance importante ce qui fait qu'on
n'arrive pas à trouver une coupure net pour la reparer.

Les équipes sont en train de regarder comment reparer au plus la panne.
Soit une solution provisoire: sur 200m passer une fibre à l'exterieur
et faire les soudures (un peu comme l'autre opérateur qui a eu le
probleme avec nous), soit remplacer de suite les 2 km de fibres (car
au final il faudra bien remplacer le bout abimé et c'est de suite
2km). La fibre on l'a. C'est pas le probleme. On regarde ce qui est
plus rapide compte tenu du type de probleme: on remplace 100m ? 500m ?
1km ? 2km ? car si on remplace que 100m c'est pas sûr que tout sera
reparé et plus la distance est grande plus de temps ça va prendre.

Il n'y a pas de delai de remise en service.


Comment by OVH - Saturday, 08 September 2007, 00:34AM

better news. We succed in pulling the first 650 m cable. We are going on to repa
ir the 48 FO - Estimated end time : 9:00 PM GMT.
In an other en we are progressing slowly as it is a small forest truck to the th
eorical manhole where we identified the second fault. As soon as we find it, we
will begin to dig. After we will have to pull from the first impact point 1383 m
cable.
We think, except new difficulties, we could finish something around 7:00 am GMT
to-morrow morning.


Next update arround 8:00 PM GMT

----------------


Please be informed that since 06:00 GMT there are 6 engineers and 3 civilian wor
ker teams looking forward that big trouble.

We go on reparing the first 640 m cable. Expected end time around 09:00 PM GMT.
At the moment we dig out one more manhole (NUM.CT.MPE). We are diging on a 10 m
square around the theorical manhole position in the forest at 1383 meters from t
he first impact.
The others civilian workers are preparing the excavation to pull the cable under
3 roads.
If everything ok, we hope to repair that cable to-morrow morning around 7:00 am
GMT.


Next update will be done around 10:00 PM GMT by our call center.


Comment by OVH - Saturday, 08 September 2007, 10:27AM

Just to keep you informed, the first chamber at 650 meters is connected now.
The second chamber has been found and they started to pull the new cable on
1383 meters.

Next update will be done by our call-Center, when the cable will be to pull,
aro und 03:00 AM GMT.
I think that now the end of intervention will be around 10:00 AM GMT.


Comment by OVH - Saturday, 08 September 2007, 10:28AM

We just proceeded to change over 2 engineers.
We waste a lot of time to find the last manhole in the forest as the drawings
are false.
We got difficulties in pulling the cable for 400 meters. It was blocked in the
sheath. We have to dig at that point and to put the cable directly on the earth.

At the moment :
they pulled the cable for 1000 m. One team began to prepare the cable in the 1st
manhole.
We are keeping on to pull the cable for the last 400 m.
I think the first 4 active FO will be connected around 9:00 am GMT.

Next update = around 07:00 am


Comment by OVH - Saturday, 08 September 2007, 15:53PM

à 14h07 la fibre a été réparée et le lien est à nouveau up.

Un sacré travail de terrain !


Comment by OVH - Saturday, 08 September 2007, 19:36PM

Les travaux continuent. Une microcoupure à 16h11


Comment by OVH - Saturday, 08 September 2007, 19:38PM

PEC ?AC1 fibre link still down. OTDR measurements now in prog
TI yet to confirm DF status.

---

PEC-AC1 Link still down. Team to send eng?s to Paris Trembley to carry out
OTDR measurements.
TI still unable to confirm DF status.

---

Eng?s now in both Paris and Trembley to migrate PEC-AC1 to spare fibre and carry
out OTDR measurements

---

PEC AC1 link restored. All alarms cleared. Span and PM check in prog

---

PEC AC1 moved from fp 3,4 to 5,6.
PM and span for PEC AC1 ok.
TI DF also migrated from 29,30 to 31,32.
TI confirmed link ok.
OTDR measument showed another cut about 1 km from the last repair point.

NB: This is only a temporary fix. Team to raise PA for permanent fix.


Comment by OVH - Saturday, 08 September 2007, 19:41PM

Parent Case Change Status from status "Tech in Route" to status "Monitoring"
Status: 'Monitoring'.


Comment by OVH - Saturday, 08 September 2007, 23:36PM

18h02 des pertes et les erreurs sur le lien pendant quelques secondes.